OpenStudy (anonymous):

what is the answer to x squared plus four x minus 5

OpenStudy (anonymous):

You would want to factor it out so we are looking for two numbers that multiply to 4 and add up to -5

OpenStudy (anonymous):

x^2+5x-x-5=0 x(x+5)-1(x+5)=0 (x-1)(x+5)=0 x=1;x=-5

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

okay so the correct answer would be (x+1) (x-5) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

unless you have to break it down even more and then you have x = 1 and x = -5

OpenStudy (anonymous):

I would put both :)